OpenStudy (anonymous):

Where will the projectile y=(vSin(A))(x/vCos(A)) - .5g(x/vCos(A))^2 hit the ground. I know that Y would be 0 but I have no idea what X would be.

OpenStudy (anonymous):

y=(vSin(A))(x/vCos(A)) - .5g(x/vCos(A))^2 0=(vSin(A))(x/vCos(A)) - .5g(x/vCos(A))^2 (vSin(A))(x/vCos(A)) =.5g(x/vCos(A))^2 vSin(A) =.5g(x/vCos(A)) can u solve for x now?

OpenStudy (anonymous):

Sure can, I finally understand it now. Thanks for the help!

OpenStudy (anonymous):

You're welcome! :)